4^x-2^x=0
slove the equation.

Answer by venugopalramana(3286)

4^x-2^x=0
slove the equation.
put 2^x=y
4^x = [2^2]^x=2^(2x)=[2^x]^2=y^2
y^2-y=0
y(y-1)=0
y=0.....or......y=1
2^x=0....not possible
2^x=1
x=0